A spherical balloon of 32 cm diameter is to be filled up with He at NTP from a cylinder containing the gas at 23 atm at 27°C. The cylinder can hold 5.42 litre of He at NTP. Calculate the number of balloons that can be filled up. 

Grade:11

1 Answers

Vikas TU
5 years ago
Number of moles the balloon can withhold from PV = nRT
n = PV/RT => (23*0.0171)/(0.0821*300) => 0.01596833
 
Number of moles cylinder can fill = 5.42/22.4 = > 0.2419
 
Number of balloons can be filled = > 0.2419/0.0159683 = 16 approx.